OpenStudy (anonymous):

The widths of two similar rectangles are 16 cm and 14 cm. What is the ratio of the perimeters? Of the areas? A) 8 : 7 and 64 : 49 B) 9 : 8 and 64 : 49 C) 9 : 8 and 81 : 64 D) 8 : 7 and 81 : 64 Thanks!

OpenStudy (anonymous):

16/14 = 8/7 For the areas: 8*8=64, 7*7 = 49 A, my best guess!